• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name johna

    Log entry time 00:26:41 on March25,2005

    Entry number 142343

    keyword=Shift summary

    Start of shift: No beam - power supply in "north arc" is being fixed

    19:30 according to ELOG, power supply is fixed, waiting for PSS certification

    20:45 MCC tells us that the power supply has been fixed, and they'll try to deliver beam in ~10 minutes.

    21:30 Get beam, ask for 5uA to do a harp scan. Beam is roughly 110 by 230 um. Ask for 40 uA and start data taking.

    Took three good runs: 2914, 2915, 2916. 2916 stopped due to HV trips. The HV gui was not responding, Steve Wood is busy trying to restore communication and high voltage at shift's end.